OpenStudy (anonymous):

Write the log equation and then solve using logs: x = [(72.43^(1/5))(.00035^(2/3))] / 5.84^(1/6)

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

\[\text{Log10}\left[\frac{72.43^{1/5} \left(0.00035^2\right)^{1/3}}{5.84^{1/6}}\right]=-2.05971 \]\[10^{-2.05971}=0.00871545 \]Computed directly by Mathematica:\[(72.43{}^{\wedge}(1/5)*0.00035{}^{\wedge}(2/3))/5.84{}^{\wedge}(1/6)=0.00871553 \]